# Hydrodynamic behavior near dynamical criticality of a facilitated conservative lattice gas

### v1

@@ Abstract
Models displaying dynamical phase transitions have attracted increasing scrutiny in recent years. Such models are tightly related to self-organizing criticality: open systems dynamically adjust their density in order to reach a critical state at density $\rho_{c}$, often displaying non-trivial scaling properties. This phenomenon manifests itself, in a closed system, as a dynamical phase transition: below $\rho_{c}$, the system reaches an absorbing state, while above $\rho_{c}$, it remains in a quasi-stationary active state.
A fundamental example of such a model is the constrained conservative lattice gas (CLG) [1], also referred to as facilitated exclusion process in the recent literature. It is defined as an exclusion particle system (i.e. any system site cannot contain more than one particle) on a $d$-dimensional lattice, where so-called active particles jump randomly at rate $1$ to each empty nearest neighbor [^1] Sometimes, simultaneous jumps are considered [1], but this does not change the critical behavior of the model.. A particle is considered active if at least one of its neighboring sites is also occupied, and the total number of particles is conserved. Related models featuring an absorbing phase transition have generated an intense research activity, like for instance the paradigm Manna sandpile model [3, 4, 5, 6]. In particular many of these models, including the CLG, exhibit a hyperuniform critical state [7, 8], for which we still have a limited knowledge. The CLG has been investigated numerically in [1, 9, 7, 8] when $d\geq 2$ and theoretically in [10, 11, 12, 13, 14, 15, 16] when $d=1$. We focus here on the $2$-dimensional case, and recall some previous results already obtained in the $1$-dimensional case.
@@
Figure 1: Median absorption time in a closed box, of size changing from 5 to 100. In the subcritical phase (a) the absorption time grows sub-linearly with the system size. In the supercritical phase (b) it grows exponentially fast, and the figure illustrates nicely that the geometric correlation length, which separates the absorbing regime from the quasi-stationary regime, is roughly $\xi\approx 20$ at $\rho=0.334$.
@@
  • The geometric correlation length $\xi_{\perp}$: this is the scale which is mostly used in the literature [18, Section 3.3], and is the one discussed in [1]. It describes the spread of activity. More precisely, to sustain activity, particle clusters must self-activate, i.e., a particle activates its neighbor, which further activates its neighbors, until closing a cycle and re-activating the particle we started with. The diameter of this self-activated structure is described by the geometric correlation length $\xi_{\perp}$.
@@
We are interested in the hydrodynamic behavior of the CLG, whose diffusion coefficient $D(\rho)$ behaves, close to $\rho_{c}$, as $(\rho-\rho_{c})^{\alpha}$ for some exponent $\alpha$. In order to understand the noise’s amplitude, we also consider the compressibility $\chi(\rho)\sim(\rho-\rho_{c})^{\gamma}$. For the CLG, the dominant parameter for the system is the density of active particles, $\rho_{a}(\rho)$. However, the notion of active particles is in fact ambiguous, since one may or may not count as active particles who are fully surrounded by other particles (and therefore cannot move). For this reason, we distinguish between $\rho_{a}(\rho)\sim(\rho-\rho_{c})^{\beta}$, the density of particles having at least one occupied neighbor (which is the one considered in [1]), and the activity $a(\rho)\sim(\rho-\rho_{c})^{b}\leq 3\rho_{a}$, which counts the local number of possible jumps. We will see further that in fact both exponents $\beta$ and $b$ coincide (Fig. 3). This means that the perimeter of clusters of active particles is of the same order as their volume.
@@
This relation can be obtained by considering the structure factor $S_{\rho}(k)$ (see [22, Section II.2.1]), which encapsulates the $2$-point statistics of the distribution at a fixed time [19]. By the scaling hypothesis, $S_{\rho}(k)$ can only depend on $k$ via the combination $\xi_{\times}k$. Moreover, $S_{\rho}(0)=\chi(\rho)$, and at criticality $S_{\rho_{c}}(k)\sim C|k|^{d-2\zeta}$ when $|k|$ is small. These three facts impose the form
@@
The next relation stems from Einstein’s relation $D=\sigma/\chi$ and the fact that the noise amplitude is determined by the number of possible particle jumps $\sigma=a$ (see [22, Section II.2] for instance): this leads to
@@
Hydrodynamics and scale invariance. Near criticality, we are interested in the macroscopic evolution of $u:=\rho-\rho_{c}$. It evolves according to the fluctuating hydrodynamic equation
@@
We emphasize that on a length above $\xi_{\times}$ the scale invariance is not the same, and in particular the dynamic exponent $z$ will change. This scale separation has been noted qualitatively in [1].
@@
Table 1: Critical exponents related to observables, in $d=1$ and $d=2$. The $1$-dimensional exponents are exact, see below. In the $2$-dimensional case, the first line is obtained either directly from our simulation results, or extracted from scaling relations. The second line contains simulation results taken from previous articles. In our case, the exponents $\nu_{\times}$ and $\zeta$ are extracted from the simulation of Fig. 4. The exponents $\beta$ and $b$ are taken from the one of Fig. 3a. The exponent $\gamma$ is simulated twice: the value $1$ obtained from Fig. 3 and equation (4). The value $1.07$ is obtained from Fig. 4. The exponent $\alpha$ is calculated using equation (7), $\nu_{\perp}$ from equation (3) and $z$ from (11).
@@
and critical exponent $\alpha=0$. The explicit construction of the grand-canonical state $\pi_{\rho}$ yields the other observables for $\rho\geq\rho_{c}$, as well as their critical exponent (see [15]): namely the activity $a(\rho)=\rho^{-1}(1-\rho)(2\rho-1)$ with $b=1$, and the compressibility $\chi(\rho)=\rho(1-\rho)(2\rho-1)$, with $\gamma=1.$ Moreover, the stationary measure can be seen as a nearest-neighbor spin system with chemical potential $\mu$, and an interaction which gives infinite costs to two neighboring empty sites. This can be solved using standard methods involving the transfer matrix (see [24, Chapter 6]), which here is given by
@@
Numerical simulations. We note that in finite systems the critical density depends slightly on the geometry, so that in the analysis of the simulated data we do not enforce a single critical density for systems of different sizes or boundary conditions.
@@
Figure 4: We show here a collapse of $S_{\rho}(k)$ for different values of $\rho$. Fitting the data shows that $\chi(\rho)=0.61(\rho-0.3361)^{-1.07}$ and $\xi_{\times}=0.03(\rho-0.3361)^{-1.77}$. Indeed, after this rescaling the curves $S_{\rho}(k)$ collapse as expressed in equation (2). The fit (black curve) is given by $S_{\rho}(k)=0.61(\rho-0.3361)^{1.07}+0.07|k|^{0.60}$.
@@
Conclusion. In this article, we discussed the critical scaling for the CLG. We saw that there are three independent critical exponents, $\beta$, $b$, and $\zeta$, that all other exponents ($\alpha,\gamma,\nu_{\perp},\nu_{\times},z$) could be deduced from. Due to repulsion, active cluster sizes are of order $1$, so their perimeter is proportional to the volume, thus $\beta=b$ and all scaling is described by two independent exponents.
We numerically computed several critical exponents for the $2d$–CLG (see Table 1), and confronted them both with those critical relations, and the numerical values in [1] and [7] for CLG with simultaneous jumps. We obtained very good agreement between them; with the exception of $\zeta$ and $\nu_{\times}$. While our numerical values fit the theoretical relations introduced above, they are different from those of [7]. The reason seems to be that we approach the critical state from $\rho>\rho_{c}$, while [7] do from $\rho<\rho_{c}$. Recently, [8] went further investigating the approach to hyperuniformity from the subcritical regime; on the contrary to [7], they find that the critical exponent $\nu_{\times}$ (which they denote $\gamma_{1}$) is different from $\nu_{\perp}$.

### v2

@@ Abstract
Models displaying dynamical phase transitions have attracted increasing scrutiny in recent years. Such models are tightly related to “self-organized criticality”, and may also illustrate how hyperuniform structures [1] emerge in nature [2, 3]. Their complexity prevents from building a universal framework, this is why some paradigmatic models are currently under deep mathematical and physical investigation (as for instance sandpiles [4], or random organization models [5]). Open systems dynamically adjust their density in order to reach a critical state at density $\rho_{c}$, often displaying non-trivial scaling properties. This phenomenon manifests itself, in a closed system, as a dynamical phase transition: below $\rho_{c}$, the system reaches an absorbing state, while above $\rho_{c}$, it remains in a quasi-stationary active state.
A fundamental example of such a model is the constrained conservative lattice gas (CLG) [6], also referred to as facilitated exclusion process in the recent literature. It is defined as an exclusion particle system (i.e. any system site cannot contain more than one particle) on a $d$-dimensional lattice, where so-called active particles jump randomly at rate $1$ to each empty nearest neighbor  [^1] Sometimes, simultaneous jumps are considered [6], but this does not change the critical behavior of the model.. A particle is considered active if at least one of its neighboring sites is also occupied, and the total number of particles is conserved. Related models featuring an absorbing phase transition have generated an intense research activity, like for instance the paradigm Manna sandpile model [8, 9, 10, 11]. In particular many of these models, including the CLG, exhibit a hyperuniform critical state [12, 13], for which we still have a limited knowledge. The CLG has been investigated numerically in [6, 14, 12, 13] when $d\geqslant 2$ and theoretically in [15, 16, 17, 18, 19, 20, 21] when $d=1$. We focus here on the $2$-dimensional case, and recall some previous results already obtained in the $1$-dimensional case. While we expect to see the same general picture in higher dimensions, numerical studies of the model become more complicated. Some critical exponents in $d=3$ are found in [12], and we expect the critical relations laid out in this paper to hold (though we do not provide verification).
XXX
Figure 1: Blue circle particles are active, red square particles are frozen. As an example, the active particle highlighted with $\square$ can jump to one of its three neighbours indicated with X. The dashed region corresponds to the frozen phase.
@@
Figure 2: Median absorption time in a closed box, of size changing from 5 to 100. In the subcritical phase (a) the absorption time grows sub-linearly with the system size. In the supercritical phase (b) it grows exponentially fast at large $L$. In fact, the figure illustrates nicely that at $\rho=0.334$ the geometric correlation length, which separates the absorbing regime from the quasi-stationary regime, is roughly $\xi\approx 20$, which is the approximative point where the exponential growth begins.
@@
  • The geometric correlation length $\xi_{\perp}$: this is the scale which is mostly used in the literature [23, Section 3.3], and is the one discussed in [6]. It describes the spread of activity. More precisely, to sustain activity, particle clusters must self-activate, i.e., a particle activates its neighbor, which further activates its neighbors, until closing a cycle and re-activating the particle we started with. The diameter of this self-activated structure is described by the geometric correlation length $\xi_{\perp}$.
@@
We are interested in the hydrodynamic behavior of the CLG, whose diffusion coefficient $D(\rho)$ behaves, close to $\rho_{c}$, as $(\rho-\rho_{c})^{\alpha}$ for some exponent $\alpha$. In order to understand the noise’s amplitude, we also consider the compressibility $\chi(\rho)\sim(\rho-\rho_{c})^{\gamma}$, defined as the sum of the two-point correlation function over the infinite lattice. For the CLG, the dominant parameter for the system is the density of active particles, $\rho_{a}(\rho)$. However, the notion of active particles is in fact ambiguous, since one may or may not count as active particles who are fully surrounded by other particles (and therefore cannot move). For this reason, we distinguish between $\rho_{a}(\rho)\sim(\rho-\rho_{c})^{\beta}$, the density of particles having at least one occupied neighbor (which is the one considered in [6]), and the activity $a(\rho)\sim(\rho-\rho_{c})^{b}\leq 3\rho_{a}$, which counts the local number of possible jumps. We will see further that in fact both exponents $\beta$ and $b$ coincide (Fig. 4). This means that the perimeter of clusters of active particles is of the same order as their volume.
@@
This relation can be obtained by considering the structure factor $S_{\rho}(k)$ (see [26, Section II.2.1]), which encapsulates the $2$-point statistics of the distribution at a fixed time [24]. By the scaling hypothesis, $S_{\rho}(k)$ can only depend on $k$ via the combination $\xi_{\times}k$. Moreover, $S_{\rho}(0)=\chi(\rho)$, and at criticality $S_{\rho_{c}}(k)\sim C|k|^{d-2\zeta}$ when $|k|$ is small. These three facts impose the form
@@
The next relation stems from Einstein’s relation $D=\sigma/\chi$ (see [26, (2.72), Section II.2.5]) and the fact that the noise amplitude is determined by the number of possible particle jumps $\sigma=a$ (see [26, Section II.2] for instance): this leads to
@@
Hydrodynamics and scale invariance. Near criticality, we are interested in the macroscopic evolution of $u:=\rho-\rho_{c}$. It evolves according to the fluctuating hydrodynamic equation (e.g. [26, II.2.9])
@@
We emphasize that on a length above $\xi_{\times}$ the scale invariance is not the same, and in particular the dynamic exponent $z$ will change (see [25]). This scale separation has been noted qualitatively in [6].
@@
Table 1: Critical exponents related to observables, in $d=1$ and $d=2$. The $1$-dimensional exponents are exact, see below. In the $2$-dimensional case, the first line is obtained either directly from our simulation results, or extracted from scaling relations. The second line contains simulation results taken from previous articles. In our case, the exponents $\beta$ and $b$ are taken from the simulation of Fig. 4a. The exponent $\gamma$ is simulated twice: the value $1$ obtained from Fig. 4 and equation (4) ; the value $1.07$ is obtained from Fig. 5. The exponent $\nu_{\times}$ is extracted from the simulation of Fig. 5 and then $\zeta$ is computed using (1) with $\gamma=1$. The exponent $\alpha$ is calculated using equation (7), $\nu_{\perp}$ from equation (3) and $z$ from (11).
@@
and critical exponent $\alpha=0$. The explicit construction of the grand-canonical state $\pi_{\rho}$ yields the other observables for $\rho\geq\rho_{c}$, as well as their critical exponent (see [20]): namely the activity $a(\rho)=\rho^{-1}(1-\rho)(2\rho-1)$ with $b=1$, and the compressibility $\chi(\rho)=\rho(1-\rho)(2\rho-1)$, with $\gamma=1.$ Moreover, the stationary measure can be seen as a nearest-neighbor spin system with chemical potential $\mu$, and an interaction which gives infinite costs to two neighboring empty sites. This can be solved using standard methods involving the transfer matrix (see [28, Chapter 6]), which here is given by
@@
Numerical simulations. We note that in finite systems the critical density depends slightly on the geometry, so that in the analysis of the simulated data we do not enforce a single critical density for systems of different sizes or boundary conditions. Rather, we leave $\rho_{c}$ as a parameter for the regression. Since $L$ equals $300$ in one simulation and $100$ in the other, it is not surprising that we obtain values of $\rho_{c}$ who differ by $\mathcal{O}(1/L)$.
@@
Figure 5: We show here a collapse of $S_{\rho}(k)$ for different values of $\rho$. The parameters $C_{1}=0.03,C_{2}=0.71,\rho_{c}=0.3361,\nu_{\times}=1.77,\gamma=1.07$ are adjusted to best fit equation (2). That is, we find that $\chi(\rho)=0.61(\rho-0.3361)^{-1.07}$ and $\xi_{\times}=0.03(\rho-0.3361)^{-1.77}$. Indeed, after this rescaling the curves $S_{\rho}(k)$ collapse as expressed in equation (2). The fit (black curve) is given by $S_{\rho}(k)=0.61(\rho-0.3361)^{1.07}+0.07|k|^{0.60}$.
@@
Conclusion. In this article, we discussed the critical scaling for the CLG. We saw that there are three independent critical exponents, $\beta$, $b$, and $\zeta$, that all other exponents ($\alpha,\gamma,\nu_{\perp},\nu_{\times},z$) could be deduced from. Moreover, due to repulsion, active cluster sizes are of order $1$, so their perimeter is proportional to the volume, thus $\beta=b$ and all scaling is described by two independent exponents. In fact, we expect the scaling relations stated here to hold in a much larger generality than the $2$-dimensional CLG, and that $\gamma=1$ should hold as well in repulsive 2D systems, see [22] for similar relations and comparison to other models. Note that we have used the gradient condition in order to derive equations (4) and (7). This a condition is very sensitive to small changes in the dynamics, but we believe that scaling exponents and relations are universal, and depend much less on perturbations of the dynamics. At the same time, it is worth noting that some “highly non-gradient” systems are known not to satisfy these relations (e.g.  the Kob-Andersen model, in which $\alpha=\infty$ [29, 30]), but this phenomenon is due to the formation of very specific cooperative structures.
We numerically computed several critical exponents for the $2d$–CLG (see Table 1), and confronted them both with those critical relations, and the numerical values in [6] and [12] for CLG with simultaneous jumps. We obtained very good agreement between them; with the exception of $\zeta$ and $\nu_{\times}$. While our numerical values fit the theoretical relations introduced above, they are different from those of [12]. The reason seems to be that we approach the critical state from $\rho>\rho_{c}$, while [12] do from $\rho<\rho_{c}$. Recently, [31, 13] went further investigating the approach to hyperuniformity from the subcritical regime; on the contrary to [12], they find that the critical exponent $\nu_{\times}$ (which [13] denote $\gamma_{1}$) is different from $\nu_{\perp}$. That is, a separation between two different length scales is also present in the subcritical regime, but with an exponent $\nu_{\times}$ different from the supercritical one.
@@ Acknowledgements.
A.S., C.E. and M.S. contributed equally to this work; A.R. conducted the numerical simulations.